2x + 5y =-6
4x +y =-12

Answer by stanbon(48546)

2x + 5y =-6
4x +y =-12
------------------
Multiply the 1st 2 to get:
4x + 10y = -12
4x + y = -12
-------------------
Subtract 2nd from 1st to get:
9y = 0
y = 0
-----------
If y = 0 solve for "x" in 2x+5y = -6
2x + 5*0 = -6
2x = -3
x = -3
-----------
Cheers,
Stan H.